### v2

# Explainable and Class-Revealing Signal Feature Extraction via Scattering Transform and Constrained Zeroth-Order Optimization

Abstract
We propose a new method to extract discriminant and explainable features from a particular machine learning model, i.e., a combination of the scattering transform and the multiclass logistic regression. Although this model is well-known for its ability to learn various signal classes with high classification rate, it remains elusive to understand why it can generate such successful classification, mainly due to the nonlinearity of the scattering transform. In order to uncover the meaning of the scattering transform coefficients selected by the multiclass logistic regression (with the Lasso penalty), we adopt zeroth-order optimization algorithms to search an input pattern that maximizes the class probability of a class of interest given the learned model. In order to do so, it turns out that imposing sparsity and smoothness of input patterns is important. We demonstrate the effectiveness of our proposed method using a couple of synthetic time-series classification problems.
Index Terms:
Nonlinear discriminant feature extraction, scattering transform, wavelets, zeroth-order optimization, sparsity and smoothness of signals
I Introduction
In signal and image classifications, the Scattering Transform (ST) [1, 2], which cascades wavelet transform convolutions with modulus nonlinearities (i.e., absolute values) and averaging operators, has emerged as an alternative to the popular Convolutional Neural Networks (CNNs)/Deep Learning (DL) [3, 4]. Since only two or three layers of the cascades in the ST are sufficient and since it uses the predefined wavelet convolution filters, it has a number of advantages over CNNs/DL for signal classification problems: 1) its training process is computationally faster; 2) it does not require a large number of training samples; 3) it automatically generates multiscale/multifrequency feature representations of input data in an explicit manner via wavelet filters; 4) the computed ST coefficients can be fed to any classifier of choice, e.g., Multiclass Logistic Regression (MLR), Support Vector Machine (SVM), $k$-Nearest Neighbors ($k$NN), etc.; and 5) it is more mathematically tenable. Yet, it still extracts robust and quasi-invariant features of input signals relative to certain types of deformations or perturbations (e.g., a small amount of shifts, warps, noise, etc.) so that it provides comparable classification performance as DL models. In fact, for a small number of training samples, it outperforms DL models [5, 2, 6].
Our earlier work has obtained excellent results using the ST and its variant in a number of signal classification problems including: underwater object classification from acoustic wavefields [7, 8]; texture image classification [9]; and music genre classification [10]. However, one critical thing remains to be understood: Why does this combination of the features and the classifier work so well? What are the features in the original time domain that contributed to these excellent classification results? Because the ST is a nonlinear transform like DL, even if we can identify which ST coefficients significantly contributed to the correct classification, the true meaning of such coefficients in the original time domain has been elusive: the only explicitly available information is the so-called path information, i.e., a set of indices of the wavelet filters (e.g., passband info) used in all the previous layers and the current layer in order to generate that particular ST coefficient.
Hence, our aim here is to extract explainable or class-revealing features using the logistic regression coefficients learned on the ST coefficients of given training samples. More specifically, we plan to estimate signals that maximize the class probability using their ST coefficients and the trained MLR classifier for a signal class of interest. If we are successful, then we can deepen our understanding of the features in the original time domain that are responsible for discriminating one class from the other. Such understanding and insight may uncover the secret of physical nature of a given problem, e.g., detection and classification of underwater objects via scattered acoustic wavefields; see [7, 8] for the background information and our previous effort. This is quite important since it may provide us with an opportunity to design new transmitter signal patterns and/or sensors that specifically focus on discriminating a particular class of objects.
II Our Proposed Method
To be concrete, let us consider a typical signal classification problem. Let $\{\boldsymbol{x}_{i}\}_{i=1:N}$ be available training signals each of which has ${d}$ time samples. Each $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ has a class label, $k\in\{1,\ldots,K\}$, i.e., $\boldsymbol{x}_{i}\in X_{k}$, where $X_{k}\subset\mathbb{R}^{d}$ is a space of class $k$ signals. Our proposed procedure is the following.
Step 1:
Apply the ST to the training samples;
Step 2:
Train the GLMNet classifier (= the MLR classifier with the Lasso penalty [11]), which can efficiently select a small number of the ST coefficients as key features; let $(\alpha_{k},\boldsymbol{{\beta}}_{k})$, $k=1:K$ be the resulting intercepts and regression coefficient vectors;
Step 3:
Find an input pattern $\hat{\!\boldsymbol{x}}\in\mathbb{R}^{d}$ for class $k$ that minimizes the following criterion:
$\hat{\!\boldsymbol{x}}=\arg\min_{\boldsymbol{x}\in\mathbb{R}^{d}}\frac{1}{p_{k}(\boldsymbol{x})}+\mu\|\boldsymbol{x}\|_{1}+\nu\|\nabla\boldsymbol{x}\|_{2},$ (1)
where $p_{k}(\boldsymbol{x})\,:=\,\exp(\alpha_{k}+\boldsymbol{{\beta}}_{k}\cdot{\mathcal{S}}[\boldsymbol{x}])/\sum_{j=1}^{K}\exp(\alpha_{j}+\boldsymbol{{\beta}}_{j}\cdot{\mathcal{S}}[\boldsymbol{x}])$ is the probability of a signal $\boldsymbol{x}$ belonging to class $k$ (according to the trained GLMNet classifier); ${\mathcal{S}}[\boldsymbol{x}]$ is the ST coefficient vector computed from $\boldsymbol{x}$; and $\mu>0$, $\nu>0$ are the Lagrange multiplier parameters to be adjusted. Note that the minimization of the first term $1/p_{k}(\boldsymbol{x})$ is clearly equivalent to the maximization of $p_{k}(\boldsymbol{x})$ while the second and third terms promote sparsity and smoothness of $\boldsymbol{x}$, respectively.
Perhaps, it is a good idea to describe the ST a bit more precisely here. Since we only consider the second-layer ST coefficients, we define it for a given input 1D signal $\boldsymbol{x}\in\mathbb{R}^{d}$ as follows:
${\mathcal{S}}[\boldsymbol{x}]\,:=\,\left\{\rho\left(\rho\left(\boldsymbol{x}\circledast_{r_{1}}\boldsymbol{{\psi}}^{(1)}_{\lambda_{1}}\right)\circledast_{r_{2}}\boldsymbol{{\psi}}^{(2)}_{\lambda_{2}}\right)\circledast_{r_{a}}\boldsymbol{{\phi}}_{J}\right\}_{\lambda_{1}\in\Lambda_{1};\lambda_{2}\in\Lambda_{2}},$ (2)
where $\rho$ is the elementwise modulus operator, $\rho:\boldsymbol{y}=(y_{1},\ldots,y_{d})^{\scriptscriptstyle{\mathsf{T}}}\in\mathbb{R}^{d}\mapsto|\boldsymbol{y}|\,:=\,(|y_{1}|,\ldots,|y_{d}|)^{\scriptscriptstyle{\mathsf{T}}}\in\mathbb{R}_{\geq 0}^{d}$; $\circledast_{r}$ indicates a 1D circular convolution (typically done via multiplication in the Fourier domain) followed by subsampling with rate $r\geq 1$; $\boldsymbol{{\psi}}^{(l)}_{\lambda_{l}}$, is the mother wavelet filter with frequency-band index $\lambda_{l}\in\Lambda_{l}$ for the $l$th layer, $l=1,2$; and finally $\boldsymbol{{\phi}}_{J}$ is the father wavelet (i.e., lowpass) filter at scale $J\in\mathbb{N}$. Note that we allow different subsampling rates for each layer, i.e., $r_{1},r_{2}$ and for the final averaging process $r_{a}$. For the details, see [1, 6, 12] as well as [8, Chap. 3].
Also, we should explain Step 3 above a bit more here. To solve the minimization problem Eq. (1), we use the so-called zeroth-order (ZO) or derivative-free optimization [13, 14, 15]. For our proposed method, it is better to use ZO optimization algorithms than the popular first-order (FO) optimization algorithms that require computing the gradient of the objective function to be minimized. This is because the gradient $\nabla{\mathcal{S}}[\boldsymbol{x}]$ is highly discontinuous, leading update steps to converge slowly, as shown in [8, Chap. 4]. In this paper, we use the popular ZO method called Differential Evolution (DE) [16, 17, 18]. We also note that we perform the updates of solution candidates in the frequency domain (via DCT) with the pink (a.k.a. $1/f$) noise as the initial randomized candidates instead of time domain updates starting from white noise. This is because the updates of solution candidates in the time domain are too local in time and many naturally-occurring signals follow the $1/f$ power spectra [19]; see also [8, Chap. 4]. The DE is certainly not the only ZO method to use for our problems; see more discussion of the choice of ZO methods in Section IV.
III Examples
In this section, we will demonstrate the usefulness of our proposed method using two well-known synthetic signal classification problems.
III-A “Cylinder-Bell-Funnel” Signal Classification
(a) Five samples/class in the “CBF” classification problem
(b) $\boldsymbol{{\beta}}$ coefficients of “CBF” classes
(c) Solutions of Eq. (1) for the “CBF” classification problem
Figure 1: Extracted features via Eq. (1) reveal the decisive characteristics of each class in the “CBF” signal classification problem
We first tested our idea described above on a simple yet well-known time-series classification dataset, called the “Cylinder-Bell-Funnel” dataset that was introduced by one of us [20, 21] and has become quite well known in the time-series analysis literature; see, e.g., [22, 23]. In this example, we want to classify synthetic noisy signals with various shapes, amplitudes, lengths, and positions into three possible classes ($K=3$). More precisely, sample signals of the three classes were generated by:
$c(i)$ $=(6+\eta)\cdot\chi_{[a,b]}(i)+\epsilon(i)$ for “cylinder”
$b(i)$ $=(6+\eta)\cdot\chi_{[a,b]}(i)\cdot(i-a)/(b-a)+\epsilon(i)$ for “bell”
$f(i)$ $=(6+\eta)\cdot\chi_{[a,b]}(i)\cdot(b-i)/(b-a)+\epsilon(i)$ for “funnel”
where $i=1,\ldots,128$, $a$ is an integer-valued uniform random variable on the interval $[16,32]$, $b-a$ also obeys an integer-valued uniform distribution on $[32,96]$, $\eta$ and $\epsilon(i)$ are the standard normal variates, and $\chi_{[a,b]}(i)$ is the characteristic (or indicator) function on the interval $[a,b]$. Figure 1(a) shows five sample waveforms from each class. If there is no noise, we can characterize the “cylinder” signals by two step edges and constant values around the center, the “bell” signals by one ramp and one step edge in this order with positive slopes around the center, and the “funnel” signals by one step edge and one ramp in this order with negative slopes around the center. In our numerical experiments below, we used the Julia programming language [24]; more specifically, for the ST computation and the MLR with the Lasso penalty, we used our own ScatteringTransform.jl [25] and the publicly-available GLMNet.jl [26]. Then, for the DE algorithm, we used the BlackBoxOptim.jl [27] package.
We generated $100$ training signals per class and used the ST with the famous “Mexican-hat” wavelet function (i.e., the 2nd derivative of the Gaussian) as its base filter. Finally, we trained the GLMNet classifier on the 2nd layer ST coefficients $\left\{{\mathcal{S}}[\boldsymbol{x}_{i}]\in\mathbb{R}_{\geq 0}^{1078}\right\}_{i=1:300}$ of these training signals $\left\{\boldsymbol{x}_{i}\in\mathbb{R}^{128}\right\}_{i=1:300}$, to obtain the regression coefficient vector $\boldsymbol{{\beta}}_{k}\in\mathbb{R}^{1078}$ and the intercept, $\alpha_{k}\in\mathbb{R}^{1}$, $k=1,2,3$, corresponding to cylinder, bell, and funnel classes. The subsampling rates $r_{1}=r_{2}=1.5$ and $r_{a}=8$ were used while the number of wavelet filters was set as $|\Lambda_{1}|=14$ and $|\Lambda_{2}|=11$. Hence, the final output size at the 2nd layer of the ST for each input signal of length $128$ per filter pair $(\lambda_{1},\lambda_{2})$ was $\lfloor 128/1.5/1.5/8\rfloor=7$, which led to $7\times 14\times 11=1078$ ST coefficients in total. The classification accuracy of this method was almost perfect ($\approx 99$%) on newly-generated test dataset of size $3000$. Figure 1(b) displays the learned $\boldsymbol{{\beta}}_{k}$ coefficient vectors that are quite sparse thanks to the Lasso penalty while Fig. 1(c) shows the solutions of Eq. (1) using the DE algorithm. It is amazing to see that those estimated signals pinpoint the distinguishing features of those three classes: the one for the bell class and that for the funnel class are measuring the local activities around the discontinuous regions where those classes abruptly ends (bell) or starts (funnel) while it checks both edge locations for the cylinder class. They are the class-revealing features and can be viewed as the essence of the prototype signals.
III-B Triangular Waveform Classification
(a) Five samples/class of the triangular waveform problem
(b) $\boldsymbol{{\beta}}$ coefficients of the three classes
(c) Solutions of Eq. (1) for the triangular waveform problem
Figure 2: Extracted features via Eq. (1) reveal the significant characteristic of each class in the triangular signal classification problem
This is another well-known classification problem, first appeared in the classic book [28, Sec. 2.6.2]. The dimensionality (or length) of each signal was extended from the original $21$ in [28] to $128$ in order to fully utilize the power of the ST: $21$ was just too short for cascades of wavelet convolutions. Three classes of signals were generated by the following formulas:
[…]
where $i=1,\ldots,128$, $h_{1}(i)=\max(6-|i-43|/7,0)$, $h_{2}(i)=h_{1}(i-21)$, $h_{3}(i)=h_{1}(i-42)$, $u$ is a uniform random variable on the interval $(0,1)$, and $\epsilon(i)$ are the standard normal variates. Simply speaking, each class consists of random convex linear combination of two triangles with additive white Gaussian noise. Notice that the noiseless versions of these three classes form an edge of a triangular manifold in $\mathbb{R}^{128}$ where each vertex is those three triangles $\boldsymbol{h}_{k}$, $k=1,2,3$. Hence, the noisy versions are distributed within Gaussian balls around this triangular manifold. Figure 2(a) shows five sample waveforms from each class; Fig. 2(b) shows the sparse $\boldsymbol{{\beta}}_{k}$ coefficient vectors; and Fig. 2(c) displays the extracted feature/the best input pattern for each class; the red, green, and blue triangles in the background indicate the vectors $\boldsymbol{h}_{1}$, $\boldsymbol{h}_{2}$, and $\boldsymbol{h}_{3}$, respectively. We used the same parameter settings as those of the “CBF” signal classification problem. The classification rate using the ST plus MLR was $88.9$% using $1000$ newly generated signals per class as the test dataset while each of the estimated features in Fig. 2(c) gave us essentially perfect classification rate of $99.99$%. In other words, these estimated features essentially discovered the nature of class signal generation mechanism as in the “CBF” signal classification problem. The Class 1 and the Class 3 features in the bottom and top rows in Fig. 2(c) indicate the signal energy concentrates around the apices of $\{\boldsymbol{h}_{1},\boldsymbol{h}_{2}\}$ and $\{\boldsymbol{h}_{2},\boldsymbol{h}_{3}\}$, respectively while the Class 2 feature shown in the middle row of Fig. 2(c) does not have such energy concentration around $\boldsymbol{h}_{2}$, indicating that the signals of this class do not have the contribution from $\boldsymbol{h}_{2}$. This is quite interesting: if the Class 2 feature had energy concentration around $\{\boldsymbol{h}_{1},\boldsymbol{h}_{3}\}$, it would have generated worse classification rate because it would confuse some Class 1 and Class 3 signals; rather our algorithm decided to use the “nonexistence” of the $\boldsymbol{h}_{2}$ component. This has given us a serendipity: in order to understand some competing class signals, it would be better to consider feature “suppressors” instead of feature extractors!
IV Summary and Discussion
In this paper, we described our effort to understand nonlinear signal features computed by ST and the MLR classifier with the Lasso penalty with the help of the ZO optimization with constraints on sparsity and smoothness of the features in the original time domain.. Our numerical results on the two well-known synthetic signal classification problems were quite promising.
We have used the DE algorithm as our ZO method to solve Eq. (1). Clearly, other methods may work well with even faster computational speed than the DE algorithm. Hence, it is important to evaluate various ZO algorithms that is robust and computationally efficient. There are two different strategies in ZO optimization: single-particle methods and multi-particle methods. The advantage of multi-particle methods [29, 30, 31] is its population-based adaptive and evolutionary solution-search capability while the disadvantage is their difficulty to derive theoretical guarantee of convergence. Hence, we will also investigate the single-particle methods, which are theoretically more tenable and can fully utilize the advanced FO optimization algorithms once the gradient estimation is done at each iteration. In particular, we plan to investigate the so-called ZO proximal methods [32, 33] because these can handle the constraints consisting of a nonconvex and smooth function and a nonsmooth and convex function like our Eq. (1). Our preliminary numerical experiments using the PRIMA.jl [34] is quite promising: it generated similar results as those in Section III with faster computational speed than the DE algorithm.
Furthermore, we need to investigate: 1) semi-automatic determination of the optimal values of $\mu$ and $\nu$ in Eq. (1); 2) possible constraints other than sparsity and smoothness—measured by $\ell_{1}$-norm of the signal and $\ell_{2}$-norm of its gradient—are effective; and 3) the influence of the setup of the wavelet filters (e.g., type of wavelets, number of filters, their frequency overlaps, etc.) used in the ST to the estimated class-revealing features.
Finally, we want to note that our optimization strategy Eq. (1) should work in principle with any other learning model as long as it outputs the class probability $p_{k}(\boldsymbol{x})$ for a given input signal $\boldsymbol{x}\in\mathbb{R}^{d}$, e.g., a DL model equipped with “softmax” output units [4, Sec. 6.2.2.3].
Acknowledgment
This research was partially supported by the US National Science Foundation grants DMS-1912747 and CCF-1934568 as well as the US Office of Naval Research grant N00014-20-1-2381.